• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sulkosky

    Log entry time 08:00:04 on August 4,2003

    Entry number 108880

    keyword=Shift Summary

    00:00 Begin Shift with Right Q2 cycling and Septum cooling down
    00:10 Right momentum Set to 2.297 GeV/c
    00:55 MCC called to give an update. Beam will probably not be available for another two hours. But things are looking better
    02:06 MCC called, they are ready to start tuning beam. Hall A will go to Beam Permit.
    03:25 MCC called, they will need another hour to clean up the orbits
    04:15 MCC sending tuned beam to Hall A
    04:18 Right Q2 tripped off, reset worked, cycling Q2 again.
    04:24 MCC called, they will send low current CW beam.
    04:43 Septum lower coil temp at 6.85K and upper coil temp at 7.05K
    04:45 Right Q2 cycled and set. Begin ramping Septum up.
    05:00 MCC sending 5 microA of CW beam
    05:04 Set Septum currnet to 212.7A
    05:06 Ask for raster off and beam to be centered at 0,0

    Harp Scan Run number 2324, horizontal = 225 microns, vertical = 113.75 microns
    Harp Scan Run number 2325, horizontal = 237 microns, vertical = 93.92 microns
    Harp Scan Run number 2326, horizontal = 189 microns, vertical = 133.38 microns

    Run #'s 23193/3193 Raster check, 3.5 by 3.6 mm
    05:50 Move target to Pol. 3He and perform NMR
    06:12 Noticed that charge asymmetry feedback is not working. Kent paged Bryan, then restarted the Parity DAQ. Bryan cleaned up background processes.
    06:35 Ask for IHWP to be retracted
    07:00 Rotate from 180 deg to 0 deg then to 270 deg.
    07:18 Begin data taking at 270 deg. with IHWP OUT
    07:42 Ask for IHWP to be inserted
    08:00 Continue Kinematic 7.12 into the next shift.